# A Survey on Graph Neural Networks for Time Series: Forecasting, Classification, Imputation, and Anomaly Detection

### v1

@@ 4.2 Modeling Inter-Temporal Dependencies
4.3 Forecasting Architectural Fusion
@@
Discrete Architectures. Numerous existing GNN-based time series forecasting methods are discrete models. For instance, factorized approaches like STGCN [66] employ a sandwich structure of graph and temporal gated convolution layers as its fundamental building block, facilitating the modeling of inter-variable and inter-temporal relations. Subsequent works, such as DGCNN [124], LSGCN [92], STHGCN [130], and HGCN [128], retain this model architecture while introducing enhancements such as dynamic graph structure estimation [124], hypergraph convolution [128], and hierarchical graph generation [128]. A multitude of other studies adhere to similar principles, stacking diverse spatial and temporal modules in their core building blocks. For example, ST-MetaNet [82] interweaves RNN cells and GAT [117] to model evolving traffic information. Comparable works include ST-MGCN [54], DSATNET [131], and EGL [132]. In contrast, ASTGCN [84], DSTAGNN [96], and GraphSleepNet [133] are constructed upon spatial-temporal attention and convolution modules, with the latter module comprising of stacking ChebConv [108] and the convolution in the temporal dimension. Graph WaveNet [85], SLCNN [89], StemGNN [62], MTGNN [61], STFGNN [57], and TGC [28] share a similar model architecture without the attention mechanism. There are also alternative designs within the realm of discrete factorized forecasting models. For instance, STAR [93] integrates the proposed spatial and temporal Transformers, while ST-GDN [115] initially performs attention-based temporal hierarchical modeling before applying various graph domain transformations. TPGNN [97] employs temporal attention and the proposed temporal polynomial graph module to more effectively capture time-evolving patterns in time series data. MTHetGNN [59] stacks the proposed temporal, relational, and heterogeneous graph embedding modules to jointly capture spatial-temporal patterns in time series data. CausalGNN [41] models multivariate time series with causal modeling and attention-based dynamic GNN modules. Auto-STGCN [121] explores high-performance discrete combinations of different spatial and temporal modules.
@@
5 GNNs for Time Series Anomaly Detection
@@
Time series anomaly detection aims to identify data observations that do not conform with the nominal regime of the data-generating process [141]. We define anomaly as any such data point, and use the term normal data otherwise; we note however that different terminologies, like novelty and outlier, are used almost interchangeably to anomaly in the literature [142]. These deviations from the nominal conditions could take the form of a single observation (point) or a series of observations (subsequence) [143]. However, unlike normal time series data, anomalies are difficult to characterize for two main reasons. First, they are typically associated with rare events, so collecting and labeling them is often a daunting task. Secondly, establishing the full range of potential anomalous events is generally impossible, spoiling the effectiveness of supervised learning techniques. Consequently, unsupervised detection techniques have been widely explored as a practical solution to challenging real-world problems.
@@
Traditionally, methods [144] such as distance-based [145, 146, 147], and distributional techniques [148] have been widely used for detecting irregularities in time series data. The former family uses distance measures to quantify the discrepancy of observations from representative data points, while the latter looks at points of low likelihood to identify anomalies. As the data-generating process becomes more complex and the dimensionality of the multivariate time series grows these methods become less effective [149].
@@
With the advancement of deep learning, early works proposed recurrent models with reconstruction [150] and forecasting[151] strategies respectively to improve anomaly detection in multivariate time series data. The forecasting and reconstruction strategies rely on forecast and reconstruction errors as discrepancy measures between anticipated and real signals. These strategies rely on the fact that, if a model trained on normal data fails to forecast or reconstruct some data, then it is more likely that such data is associated with an anomaly. However, recurrent models [152] are found to lack explicit modeling of pairwise interdependence among variable pairs, limiting their effectiveness in detecting complex anomalies [48, 153]. Recently, GNNs have shown promising potential to address this gap by effectively capturing temporal and spatial dependencies among variable pairs [49, 70, 154].
TABLE III: Summary of representative graph neural networks for time series anomaly detection. The strategy notation: “CL”, “FC”, “RC”, and “RL” indicate “Class”, “Forecast”, “Reconstruction”, and “Relational Discrepancies”, respectively. The remaining notations are shared with Table II.
Approach Year Venue Strategy Spatial Module Temporal Module Missing Values Input Graph Learned Relations Graph Heuristics CCM-CDT [60] 2019 IEEE TNNLS RC Spatial GNN T-R No R - PC, FD MTAD-GAT [48] 2020 IEEE ICDM FC+RC Spatial GNN T-A No NR - - GDN[49] 2021 AAAI FC Spatial GNN - No NR S - GTA[154] 2021 IEEE IoT FC Spatial GNN T-H No NR S - EvoNet[155] 2021 WSDM CL Spatial GNN T-R No R - PS Event2Graph[156] 2021 arXiv RL Spatial GNN T-A No R - PS GANF[68] 2022 ICLR RC+RL Spatial GNN T-R No NR S - Grelen[157] 2022 IJCAI RC+RL Spatial GNN T-H No NR D - VGCRN[158] 2022 ICML FC+RC Spatial GNN T-R No NR S - FuSAGNet[70] 2022 KDD FC+RC Spatial GNN T-R No NR S - GTAD[159] 2022 Entropy FC+RC Spatial GNN T-C No NR - - HgAD[160] 2022 IEEE BigData FC Spatial GNN - No NR S HAD-MDGAT[161] 2022 IEEE Access FC+RC Spatial GNN T-A No NR - - STGAN[161] 2022 IEEE TNNLS RC Spatial GNN T-R No R - SP GIF [69] 2022 IEEE IJCNN RC Spatial GNN - No R - SP, PC, FD DyGraphAD[162] 2023 arXiv FC+RL Spatial GNN T-C No R - PS GraphSAD[163] 2023 arXiv CL Spatial GNN T-C No R - PS, PC
5.1 General Framework for Anomaly Detection
@@
5.2 Discrepancy Frameworks for Anomaly Detection
@@
All the proposed anomaly detection methods follow the same backbone-scorer architecture. However, the way the backbone module is trained to learn data structure from nominal data and the implementation of the scoring module differentiate these methods into three categories: Reconstruction, forecast, and relational discrepancy frameworks.
@@
Reconstruction Discrepancy. Reconstruction discrepancy frameworks rely on the assumption that reconstructed error should be low during normal periods, but high during anomalous periods. From a high-level perspective, they are fundamentally designed to replicate their inputs as outputs [167]. However, the assumption is that the backbone is sufficiently expressive to model and reconstruct well the training data distribution, but not out-of-sample data. Therefore, a reconstruction learning framework often incorporates certain constraints and regularization terms, e.g., to enforce a low-dimensional embedded code [168] or applying variational objectives [169].
@@
Once the data structure has been effectively learned, the backbone model should be able to approximate the input during non-anomalous periods, as this input would closely resemble the normal training data. In contrast, during an anomalous event, the backbone model is expected to struggle with reconstructing the input, given that the input patterns deviate from the norm and is situated outside of the manifold. Using the reconstructed outputs from the backbone module, a discrepancy score is computed by the $\textsc{Scorer}(\cdot)$ to determine whether an anomalous event has occurred. Although deep reconstruction models generally follow these principles for detecting anomalies, a key distinction between GNNs and other architectural types rests in the backbone reconstructor, $\textsc{Backbone}(\cdot)$, which is characterized by its spatiotemporal GNN implementation.
@@
MTAD-GAT [48] utilizes a variational objective [169] to train the reconstructor module. During inference, the reconstructor module will provide the likelihood of observing each input channel signal to the scorer. The scorer then summarizes the likelihoods into a single reconstruction discrepancy as an anomaly score. With the availability of reconstruction probability for each channel variable, MTAD-GAT can diagnose the anomaly scorer by computing the contribution of each variable to the discrepancy score. While MTAD-GAT shares the same variational objective as LSTM-VAE [150], MTAD-GAT differs by employing graph attention network as a spatial-temporal encoder to learn inter-variable and inter-temporal dependencies. Empirically, it is shown to outperform LSTM on the same VAE objective [^1] While MTAD-GAT also optimizes their network using forecasting objective, its ablation shows that using reconstruction alone on GNN can outperform its LSTM counterpart.. Interestingly, MTAD-GAT also shows that attention scores in the graph attention network reflect substantial differences between normal and anomalous periods.
@@
A seminal work in applying forecasting-based GNN to detect anomalies in time series data is GDN [49]. The forecaster of GDN consists of two main parts: first, a graph structure module that learns an underlying graph structure, and, second, a graph attention network that encodes the input series representation. The graph structure module computes the graph adjacency matrix for the graph attention network on the learned graph to obtain an expressive representation for making a one-step-ahead forecast. Finally, the scorer computes the forecast discrepancy as the maximum absolute forecast error among the channel variables to indicate whether an anomaly event has occurred.
Interestingly, GDN illustrates that an anomaly event may be manifested in a single variable, which acts as a symptom, while the underlying cause may be traced to a separate, root-cause variable. Hence, GDN proposed to utilize the learned relationships between variables for diagnosing the root causes of such events rather than relying solely on the individual contributions of each variable to the discrepancy score for diagnosing the root cause of anomaly events. This is accomplished by identifying the symptomatic variable that results in the maximum absolute error, followed by pinpointing its neighbour variables. The ability of GDN to discern these associations underscores the potential of GNNs in offering a more holistic solution to anomaly detection and diagnosis through the automated learning of inter-time series relationships.
@@
GReLeN [157] was the first to leverage learned dynamic graphs to detect anomalies from the perspective of relational discrepancy. To achieve this, the reconstruction module of GReLeN learns to dynamically construct graph structures that adapt at every time point based on the input time series data. The constructed graph structures serve as the inputs for a scorer that computes the total changes in the in-degree and out-degree values for channel nodes. GReLeN discovered that by focusing on sudden changes in structural relationships, or relational discrepancy, at each time point, they could construct a robust metric for the detection of anomalous events.
@@
Hybrid and Other Discrepancies. Each type of discrepancy framework often possesses unique advantages for detecting and diagnosing various kinds of anomalous events. As demonstrated in GDN[49], the relational discrepancy framework can uncover spatial anomalies that are concealed within the relational patterns between different channels. In contrast, the forecast discrepancy framework may be particularly adept at identifying temporal anomalies such as sudden spikes or seasonal inconsistencies. Therefore, a comprehensive solution would involve harnessing the full potential of spatiotemporal GNNs by computing a hybridized measure that combines multiple discrepancies as indicators for anomaly detection. For instance, MTAD-GAT[48] and FuSAGNet[70] employ both reconstruction and forecast discrepancy frameworks, while DyGraphAD[162] utilizes a combination of forecast and relational discrepancy frameworks to enhance the detection of anomalous events. In these cases, the scoring function should be designed to encapsulate the combination of either reconstruction, forecast or relational discrepancy. In general, the anomaly score can be represented as $S_{t}=\|\mathbf{X}_{t}-\mathbf{\hat{X}}_{t}\|_{2}^{2}+\|\mathbf{A}_{t}-\mathbf{\hat{A}}_{t}\|_{F}^{2}$, which captures the reconstruction and relational discrepancies, respectively. Here, $\mathbf{X}_{t}$ and $\mathbf{A}_{t}$ represent the target signals and inter-series relations, while $\mathbf{\hat{X}}_{t}$ and $\mathbf{\hat{A}}_{t}$ denote the predicted signal and inter-series relations.
@@
6 GNNs for Time Series Classification
TABLE IV: Summary of graph neural networks for time series classification. Task notation: “U” and “M” refer to univariate and multivariate time series classification tasks. Conversion represents the transformation of a time series classification task into a graph-level task as either graph or node classification task, represented as “Series2Graph” and “Series2Node”, respectively. The remaining notations are shared with Table II.
Approach Year Venue Task Conversion Spatial Module Temporal Module Missing Values Input Graph Learned Relations Graph Heuristics MTPool[175] 2021 NN M - Spatial GNN T-C No NR S - Time2Graph+[73] 2021 TKDE U Series2Graph Spatial GNN - No R - PS RainDrop[46] 2022 ICLR M - Spatial GNN T-A Yes NR S - SimTSC[74] 2022 SDM U+M Series2Node Spatial GNN T-C No R - PS LB-SimTSC[75] 2023 arXiv U+M Series2Node Spatial GNN T-C No R - PS TodyNet[176] 2023 arXiv M - Spatial GNN T-C No NR D -
@@ 6 GNNs for Time Series Classification
6.1 Univariate Time Classification
@@
In the following, we delve into two novel graph-based approaches to univariate time series classification, namely Series2Graph and Series2Node.
@@
Series2Graph. The Series2Graph approach transforms a univariate time series into a graph to identify unique patterns that enable accurate classification using a GNN. In this manner, each series is treated as a graph, and the graph will be the input for a GNN to make classification outputs.
@@
Firstly, each series is broken down into subsequences as nodes, and the nodes are connected with edges illustrating their relationships. Following this transformation, a GNN is applied to make graph classification. This procedure is represented in the upper block of 4d. Fundamentally, it seeks to model inter-temporal dependencies under a GNN framework to identify temporal patterns that differentiate series samples into their respective classes.
@@
The Series2Graph perspective was first proposed by the Time2Graph+ [187, 73] technique. The Time2Graph+ modeling process can be described as a two-step process: first, a time series is transformed into a shapelet graph, and second, a GNN is utilized to model the relations between shapelets. To construct a shapelet graph, the Time2Graph algorithm partitions each time series into successive segments. It then employs data mining techniques to assign representative shapelets to the subsequences. These shapelets serve as graph nodes. Edges between nodes are formed based on the conditional probability of a shapelet occurring after another within a time series. Consequently, each time series is converted into a graph where shapelets form nodes and transition probabilities create edges. After graph construction, Time2Graph+ utilizes Graph Attention Network along with a graph pooling operation to derive the global representation of the time series. This representation is then fed into a classifier to assign class labels to the time series.
While Time2Graph+ is the only GNN method within this framework, we use the term Series2Graph to describe the process of converting a time series classification task into a graph classification task.
@@
Series2Node. As capturing differentiating class patterns across different series data samples is important, leveraging relationships across the different series data samples in a given dataset can be beneficial for classifying a time series. To achieve this, one can take the Series2Node approach, where each series sample is seen as a separate node. These series nodes are connected with edges that represent the relationships between them, creating a large graph that provides a complete view of the entire dataset.
@@
The Series2Node was originally proposed by SimTSC [74] approach. With SimTSC, series nodes are connected using edges, which are defined by their pairwise DTW distance, to construct a graph. During the modeling process, a primary network is initially employed to encode each time series into a feature vector, thus creating a node representation. Subsequently, a standard GNN operation is implemented to derive expressive node representations, capturing the similarities between the series. These node representations are then inputted into a classifier, which assigns class labels to each time series node in the dataset. LB-SimTSC [75] extends on SimTSC to improve the DTW preprocessing efficiency by employing the widely-used lower bound for DTW, known as LB_Keogh [188]. This allows for a time complexity of $O(L)$ rather than $O(L^{2})$, dramatically reducing computation time.
@@
The Series2Node process essentially transforms a time series classification task into a node classification task. As illustrated in the lower block of 4d, the Series2Node perspective aims to leverage the relationships across different series samples for accurate time series node classification [75]. It is also an attempt to marry classical distance-based approaches with advanced GNN techniques. While not explicitly depicted in the figure, it is important to note that the same concept can be applied to classify multivariate time series by modifying the backbone network [74].
6.2 Multivariate Time Series Classification

### v3

@@ A.2 Modeling Inter-temporal Dependencies
A.3 Forecasting Architectural Fusion
@@
Discrete Architectures. Numerous existing GNN-based time series forecasting methods are models processing discrete data. For instance, factorized approaches like STGCN [54] employ a sandwich structure of graph and temporal gated convolution layers as its fundamental building block, facilitating the modeling of inter-variable and inter-temporal relations. Subsequent works, such as DGCNN [126], LSGCN [105], STHGCN [224], and HGCN [127], retain this model architecture while introducing enhancements such as dynamic graph structure estimation [126], hypergraph convolution [127], and hierarchical graph generation [127]. A multitude of other studies adhere to similar principles, stacking diverse spatial and temporal modules in their core building blocks. For example, ST-MetaNet [76] interweaves RNN cells and GAT [120] to model evolving traffic information. Comparable works include ST-MGCN [41], DSATNET [176], and EGL [225]. In contrast, ASTGCN [86], DSTAGNN [88], and GraphSleepNet [226] are constructed upon spatial-temporal attention and convolution modules, with the latter module comprising of stacking ChebConv [113] and the convolution in the temporal dimension. Graph WaveNet [48], SLCNN [95], StemGNN [50], MTGNN [51], STFGNN [44], and TGC [28] share a similar model architecture without the attention mechanism. There are also alternative designs within the realm of discrete factorized forecasting models. For instance, STAR [83] integrates the proposed spatial and temporal transformers, while ST-GDN [211] initially performs attention-based temporal hierarchical modeling before applying various graph domain transformations. TPGNN [60] employs temporal attention and the proposed temporal polynomial graph module to more effectively capture time-evolving patterns in time series data. MTHetGNN [46] stacks the proposed temporal, relational, and heterogeneous graph embedding modules to jointly capture spatial-temporal patterns in time series data. CausalGNN [227] models multivariate time series with causal modeling and attention-based dynamic GNN modules. Auto-STGCN [123] explores high-performance discrete combinations of different spatial and temporal modules.
@@
Appendix B GNNs for Time Series Anomaly Detection
@@
Time series anomaly detection aims to identify data observations that do not conform with the nominal regime of the data-generating process [134]. We define anomaly as any such data point, and use the term normal data otherwise; we note however that different terminologies, like novelty and outlier, are used almost interchangeably to anomaly in the literature [135]. Deviations from the nominal conditions could take the form of a single observation (point) or a series of observations (subsequence) [136]. However, unlike normal time series data, anomalies are difficult to characterize for two main reasons. Firstly, they are typically associated with rare events, so collecting and labeling them is often a daunting task. Secondly, establishing the full range of potential anomalous events is generally impossible, spoiling the effectiveness of supervised learning techniques. Consequently, unsupervised detection techniques have been widely explored as a practical solution to real-world problems.
@@
Traditionally, methods [233] such as distance-based [137, 234, 235], and distributional techniques [138] have been widely used for detecting irregularities in time series data. The former family uses distance measures to quantify the discrepancy of observations from representative data points, while the latter looks at points of low likelihood to identify anomalies.
@@
The advent of deep learning has sparked significant advancements, drawing lessons from earlier methods. Early research in this area proposed recurrent models with reconstruction [139] and forecasting[140] strategies respectively to improve anomaly detection in multivariate time series data. The forecasting and reconstruction strategies rely on forecast and reconstruction errors as discrepancy measures between anticipated and real signals. These strategies rely on the fact that, if a model trained on normal data fails to forecast or reconstruct some data, then it is more likely that such data is associated with an anomaly. However, recurrent models [141] are found to lack explicit modeling of pairwise interdependence among variable pairs, limiting their effectiveness in detecting complex anomalies [35, 142]. Recently, GNNs have shown promising potential to address this gap by effectively capturing temporal and spatial dependencies among variable pairs [36, 56, 89].
B.1 General Approach for Anomaly Detection
@@
B.2 Discrepancy Frameworks for Anomaly Detection
@@
Most of anomaly detection methods follow the same backbone-scorer architecture. However, the way the backbone module is trained to learn data structure from nominal data and the implementation of the scoring module differentiate these methods into three categories: (1) reconstruction, (2) forecast, and (3) relational discrepancy frameworks.
@@
Reconstruction Discrepancy. Reconstruction discrepancy frameworks rely on the assumption that reconstructed error should be low during normal periods, but high during anomalous periods. From a high-level perspective, they are fundamentally designed to replicate their inputs as outputs like autoencoders [146]. However, it is assumed the backbone is sufficiently expressive to model and reconstruct well the training data distribution, but not out-of-sample data. Therefore, a reconstruction learning framework often incorporates certain constraints and regularization terms, e.g., to enforce a low-dimensional embedded code [147] or applying variational objectives [148].
@@
Once the data structure has been effectively learned, the backbone model should be able to approximate the input during non-anomalous periods, as this input would closely resemble the normal training data. In contrast, during an anomalous event, the backbone model is expected to struggle with reconstructing the input, given that the input patterns deviate from the norm and is situated outside the manifold. Using the reconstructed outputs from the backbone module, a discrepancy score is computed by the $\textsc{Scorer}(\cdot)$ to determine whether an anomalous event has occurred. Although deep reconstruction models generally follow these principles for detecting anomalies, a key distinction between GNNs and other architectural types rests in the backbone reconstructor, $\textsc{Backbone}(\cdot)$, which is characterized by its spatiotemporal GNN implementation.
@@
MTAD-GAT [35] utilizes a variational objective [148] to train the backbone reconstructor module. During inference, the reconstructor module will provide the likelihood of observing each input channel signal to the scorer. The scorer then summarizes the likelihoods into a single reconstruction discrepancy as an anomaly score. With the availability of reconstruction probability for each channel variable, MTAD-GAT can diagnose the anomaly scorer by computing the contribution of each variable to the discrepancy score. While MTAD-GAT shares the same variational objective as LSTM-VAE [139], MTAD-GAT differs by employing graph attention network as a spatial-temporal encoder to learn inter-variable and inter-temporal dependencies. Empirically, it is shown to outperform LSTM on the same VAE objective [^1] While MTAD-GAT also optimizes their network using forecasting objective, its ablation shows that using reconstruction alone on GNN can outperform its LSTM counterpart.. Interestingly, MTAD-GAT also shows that attention scores in the graph attention network reflect substantial differences between normal and anomalous periods.
@@
A seminal work in applying forecasting-based GNN to detect anomalies in time series data is GDN [36]. The forecaster of GDN consists of two main parts: first, a graph structure module that learns an underlying graph structure, and second, a graph attention network that encodes the input series representation. The graph structure module computes the graph adjacency matrix for the graph attention network on the learned graph to obtain an expressive representation for making a one-step-ahead forecast. Finally, the scorer computes the forecast discrepancy as the maximum absolute forecast error among the channel variables to indicate whether an anomaly event has occurred.
Interestingly, GDN illustrates that an anomaly event may be manifested in a single variable, which acts as a symptom, while the underlying cause may be traced to a separate, root-cause variable. Hence, GDN proposed to utilize the learned relationships between variables for diagnosing the root causes of such events rather than relying solely on the individual contributions of each variable to the discrepancy score for diagnosing the root cause of anomaly events. This is accomplished by identifying the symptomatic variable that results in the maximum absolute error, followed by pinpointing its neighbour variables. The ability of GDN to discern these associations underscores the potential of GNNs in offering a more holistic solution to anomaly detection and diagnosis through the automated learning of inter-variable relationships.
@@
While the forecast discrepancy frameworks can be similar to reconstruction discrepancy frameworks, both rely on fundamentally different principles and thus implementation strategies. The reconstruction discrepancy framework projects the current timestamp’s input onto a learned latent space and attempts to reconstruct it, identifying anomalies when reconstruction fails, as this indicates a significant deviation from the learned normal patterns. This method directly uses current input data during both training and inference, centering on the model’s ability to reconstruct well the training data distribution but not out-of-sample data. Conversely, the forecast discrepancy framework does not directly use the current timestamp data when making a forecast. Instead, it relies on historical data to predict current values and identifies anomalies by comparing these predictions with actual current observations, flagging significant discrepancies. Thus, reconstruction focuses on accurately replicating current inputs, while forecasting emphasizes predicting future data points and detecting anomalies through prediction errors. Advanced GNN forecast discrepancy framework, GST-Pro [150], can even predict anomalies without using actual observations in the scoring function, allowing for the prediction of anomalies at future timestamps before they occur.
@@
GReLeN [90] was the first method to leverage on learned dynamic graphs to detect anomalies from the perspective of relational discrepancy. To achieve this, the reconstruction module of GReLeN learns to dynamically construct graph structures that adapt at every time point based on the input time series data. The constructed graph structures serve as the inputs for a scorer that computes the total changes in the in-degree and out-degree values for channel nodes. GReLeN discovered that by focusing on sudden changes in structural relationships, or relational discrepancy, at each time point, they could construct a robust metric for the detection of anomalous events.
@@
Hybrid and Other Discrepancies. Each type of discrepancy-based framework often possesses unique advantages for detecting and diagnosing various kinds of anomalous events. As demonstrated in GDN[36], the relational discrepancy framework can uncover spatial anomalies that are concealed within the relational patterns between different channels. In contrast, the forecast discrepancy framework may be particularly adept at identifying temporal anomalies such as sudden spikes or seasonal inconsistencies. Therefore, a comprehensive solution would involve harnessing the full potential of spatial-temporal GNNs by computing a hybridized measure that combines multiple discrepancies as indicators for anomaly detection. For instance, MTAD-GAT[35] and FuSAGNet[56] employ both reconstruction and forecast discrepancy frameworks, while DyGraphAD[82] utilizes a combination of forecast and relational discrepancy frameworks to enhance the detection of anomalous events. In these cases, the scoring function should be designed to encapsulate the combination of either reconstruction, forecast or relational discrepancy. In general, the anomaly score can be represented as $S_{t}=\|\mathbf{X}_{t}-\mathbf{\hat{X}}_{t}\|_{2}^{2}+\|\mathbf{A}_{t}-\mathbf{\hat{A}}_{t}\|_{F}^{2}$, which captures the reconstruction and relational discrepancies, respectively. Here, $\mathbf{X}_{t}$ and $\mathbf{A}_{t}$ represent the target signals and inter-variable relations, while $\mathbf{\hat{X}}_{t}$ and $\mathbf{\hat{A}}_{t}$ denote the predicted signal and inter-variable relations.
@@
Appendix C GNNs for Time Series Classification
@@
C.1 Univariate Time Series Classification
@@
In the following, we delve into two novel graph-based approaches to univariate time series classification, namely Series-as-Graph and Series-as-Node.
@@
Series-as-Graph. The Series-as-Graph approach transforms a univariate time series into a graph to identify unique patterns that enable accurate classification using a GNN. In this manner, each series is treated as a graph, and the graph will be the input for a GNN to make classification outputs.
@@
Firstly, each series is broken down into subsequences as nodes, and the nodes are connected with edges illustrating their relationships. Following this transformation, a GNN is applied to make graph classification. This procedure is represented in the upper block of Fig. 4d. Fundamentally, it seeks to model inter-temporal dependencies under a GNN framework to identify differentiating patterns across series samples into their respective classes.
@@
The graph classification perspective of Series-as-Graph was first proposed by the Time2Graph technique  [161]. This approach was later developed further with the incorporation of GNN, as Time2Graph+ [96]. The Time2Graph+ modeling process can be described as a two-step process: first, a time series is transformed into a shapelet graph, and second, a GNN is utilized to model the relations between shapelets. To construct a shapelet graph, the Time2Graph algorithm partitions each time series into successive segments. It then employs data mining techniques to assign representative shapelets to the subsequences. These shapelets serve as graph nodes. Edges between nodes are formed based on the conditional probability of a shapelet occurring after another within a time series. Consequently, each time series is transformed into a graph where shapelets form nodes and transition probabilities create edges. After graph construction, Time2Graph+ utilizes graph attention networks along with a pooling operation to derive the global representation of the time series. This representation is then fed into a classifier to assign class labels to the time series. Similarly, EC-GCN [152] constructs graphs from encrypted traffic for classification. The Series-as-Graph approach can also be extended to multivariate time-series classification task [153].
While we use Series-as-Graph to describe the formulation of a time series classification task as a graph classification task, this should not be confused with the Series2Graph [162] approach, which is utilized for the anomaly detection tasks. The strategy of recasting a series as a graph is not unique to time-series classification, as demonstrated by Series2Graph. The usage of this strategy across various tasks and domains highlights its inherent adaptability and underscores its potential worth exploring further by researchers.
@@
Series-as-Node. As capturing differentiating class patterns across different series data samples is important, leveraging relationships across the different series data samples in a given dataset can be beneficial for classifying a time series. To achieve this, one can take the Series-as-Node approach, where each series sample is seen as a separate node. These series nodes are connected with edges that represent the relationships between them, creating a large graph that provides a complete view of the entire dataset.
@@
The Series-as-Node was originally proposed by SimTSC [63] approach. With SimTSC, series nodes are connected using edges, which are defined by their pairwise DTW distance, to construct a graph. During the modeling process, a primary network is initially employed to encode each time series into a feature vector, thus creating a node representation. Subsequently, a standard GNN operation is implemented to derive expressive node representations, capturing the similarities between the series. These node representations are then inputted into a classifier, which assigns class labels to each time series node in the dataset. LB-SimTSC [97] extends on SimTSC to improve the DTW preprocessing efficiency by employing the widely-used lower bound for DTW, known as LB_Keogh [163]. This allows for a time complexity of $O(L)$ rather than $O(L^{2})$, dramatically reducing computation time.
@@
The Series-as-Node process essentially formulates a time series classification task as a node classification task. As illustrated in the lower block of Fig. 4d, the Series-as-Node perspective aims to leverage the relationships across different series samples for accurate time series node classification [97]. It is also an attempt to marry classical distance-based approaches with advanced GNN techniques. While not explicitly depicted in the figure, it is important to note that the same concept can be applied to classify multivariate time series by modifying the backbone network that transforms each series into a node [63].
C.2 Multivariate Time Series Classification